# Influence of prior and task generated emotions on XAI explanation retention and understanding

### v1

@@ Abstract
The explanation of AI results and how they are received by users is an increasingly active research field. However, there is a surprising lack of knowledge about how social factors such as emotions affect the process of explanation by a decision support system (DSS). While previous research has shown effects of emotions on DSS supported decision-making, it remains unknown in how far emotions affect cognitive processing during an explanation. In this study, we, therefore, investigated the influence of prior emotions and task-related arousal on the retention and understanding of explained feature relevance. To investigate the influence of prior emotions, we induced happiness and fear prior to the decision support interaction. Before emotion induction, user characteristics to assess their risk type were collected via a questionnaire. To identify emotional reactions to the explanations of the relevance of different features, we observed heart rate variability (HRV), facial expressions, and self-reported emotions of the explainee while observing and listening to the explanation and assessed their retention of the features as well as their influence on the outcome of the decision task. Results indicate that (1) task-unrelated prior emotions do not affected the ratantion but may affect the understanding of the relevance of certain features in the sense of an emotion-induced confirmation bias, (2) certain features related to personal attitudes yielded arousal in individual participants, (3) this arousal affected the understanding of these variables.
@@ 1 Introduction and Related Work
More recently, emotions have been investigated in the context of XAI and decision-making. In [10] it was shown that humans respond differently to explanations depending on their emotional state. Individuals with low arousal levels followed advice more when no explanation was given, whereas individuals with high arousal levels followed advice more when a guided explanation was given, i.e., an explanation that contained a context-sensitive selection of the features that were explained. It was concluded that arousal is more critical in how explanations are received than valence or the emotion category.
Furthermore, [5] have observed that negative affect can be observed when explanations are given for an easy task and positive affect in case of explanation of an AI in a difficult task, indicating that affect valence may be a useful variable to determine the explanation strategies in a specific context, i.e., whether explanations should be given. Similarly, [3] found in a vignette study that negative feelings would result from wrong advice and positive feelings from correct advice. In a further study, they found that emotions evoked by explanations increased or decreased trust [2]. Emotions together with workload can correlate with ”explanatory efficacy” [6].
[7] investigated whether XAI systems can intervene to regulate the explainees’ emotions. Here, a nudging strategy was investigated. It was found that nudging strategies to emotional debiasing are effective, yet not sufficient for rational decision-making.
However, clear guidelines regarding how emotions affect the understanding in a decision-making situation are still missing. This work contributes to the field by addressing this question.
@@ 2 Research Questions
In this research, we therefore investigate tree research questions:
@@ 3.1.1 Independent Variables
Task-related emotions. We measure the emotional reactions by using the EmoNet arousal data during the feature presentation. Based on the arousal value , an anomaly (emotional reaction) in the arousal state is detected by
@@ 3.1.3 Control Variables and Additional Measurements
mDes:
@@
EmoNet Results: In addition to arousal scores, full output vectors from the EmoNet model were stored for each facial frame, enabling detailed emotional state tracking over time.
@@ 3.2 Decision Suport System
The system centers around an embodied conversational agent named Flobi, who provides a personalized assessment of an individual’s risk profile based on a set of predefined input features. This risk assessment is subsequently used in the context of a Holt and Laury lottery task, where participants make incentivized decisions under risk. The agent’s evaluation serves as a form of decision support, offering guidance while allowing participants to ultimately make their own choices.
@@ 3.3 Participants
A total of 24 participants took part in the study. Of these, 15 were assigned to the fear condition and 9 to the happy condition. The sample included 8 male and 16 female participants. All participants were students at the Bielefeld university.
@@ 3.4 Experimental Setting — Procedure
The experiment consisted of six phases (cf. Fig. 2. We recorded the users’ heart rate variability, video and audio, and their facial expressions as computed by EmoNet. In the first phase, a questionnaire was administered that requested data from the user to assess the user’s risk type. The risk type classification was achieved by a linear scoring scheme integrating the user’s numerical answers, yielding risk a type value between 0 and 9. After an emotion induction sequence, the actual risk task was explained to the user and the user could provide his/her risk selection, i.e., selecting a high or low risk on a scale from 0 to 9. After the user’s first risk decision, the system would present its risk suggestion to the user, based on the evaluation of the user’s risk type. More specifically, a user yielding a high value for a high-risk propensity would receive a suggestion of a high-risk choice and vice versa, also on a scale from 0 to 9. This suggestion was followed by an explanation of all ten variables and their relevance to the estimated risk type of the user. For example, being female was an indicator towards less risk propensity, whereas being male was an indicator for a high-risk propensity. We analyzed the HRV and Facial Expressions during theses episodes to detect arousal. After this explanation, the user could revise his/her decision. In the last phase, the users were asked to (verbally) name the features that they remembered from the explanation. After this, they were presented the features one after another and were asked to determine whether a certain value of this feature was an indicator for higher or lower risk propensity. We used this information as a proxy for understanding.
@@ 3.4.1 Questionnaire
During the first phase of the interaction, the participants were asked to fill out an online questionnaire, with the virtual robot Floka being present and asking the questions. A total of ten questions were asked. We refer to these questions as the “variables” or “features” that the system uses to compute the risk type of the participant.
Based on results from empirical studies reported in the literature for each variable, a scoring scheme was developed that assigned a score for each answer indicating higher or lower risk propensity. Overall, this resulted in a linear scoring scheme – the more scores, the higher the risk propensity. Scores were assigned based on a median value reported in the literature above or below which a higher (+1) or lower (0) risk score was given. This resulted in a risk type value between 0 and 9 for each participant.
@@ 3.4.4 XAI decision and explanation
In the explanation phase (phase 4, see above), Flobi would then explain for each variable, whether the explainee’s (self-rated) value (e.g., of her political orientation) was an indicator adding to a higher or lower probability of the explainee as being rather more risk friendly. In this case, Flobi would say: “Because your political orientation is rather left, you are presumably more risk friendly”. Fig. 3 shows the GUI, where all variables are depicted with their respective contributions to the AI system’s decision. This would be repeated for all 11 variables.
@@ 3.4.5 Assessment of user’s retention and understanding of the explanations and the task
In the last phase the uers’ retention and understanding were assessed using the procedures described in subsection 3.1.2. See Fig. 4 for the GUI to sort the features according to the remembered influence it had on the risk type classifcation.
@@ 4.1 Emotion Manipulation Check
To examine differences in affective responses between induction conditions, we compared SAM Valence and Arousal scores across groups (Fear vs. Happy). On the Valence scale, which ranges from 1 (“unpleasant feeling”) to 5 (“pleasant feeling”), participants in the Happy condition reported higher valence ratings (median 4), indicating more pleasant emotional states compared to the Fear group (median m 3). Similarly, on the Arousal scale (1 = “calm”, 5 = “aroused”), the Fear group tended to report higher arousal levels than the Happy group, suggesting that the fear induction elicited more physiologically activating emotional responses. These trends align with theoretical expectations – fear typically evokes high arousal and negative valence, while happiness is associated with positive valence and lower arousal.

### v2

@@ Abstract
The explanation of AI results and how they are received by users is an increasingly active research field. However, there is a surprising lack of knowledge about how social factors such as emotions affect the process of explanation by a decision support system (DSS). While previous research has shown effects of emotions on DSS supported decision-making, it remains unknown in how far emotions affect cognitive processing during an explanation. In this study, we, therefore, investigated the influence of prior emotions and task-related arousal on the retention and understanding of explained feature relevance. To investigate the influence of prior emotions, we induced happiness and fear prior to the decision support interaction. Before emotion induction, user characteristics to assess their risk type were collected via a questionnaire. To identify emotional reactions to the explanations of the relevance of different features, we observed heart rate variability (HRV), facial expressions, and self-reported emotions of the explainee while they were observing and listening to the explanation and assessed their retention of the features as well as their understanding of the influence of each feature on the outcome of the decision task. Results indicate that (1) task-unrelated prior emotions do not affected the retention but may affect the understanding of the relevance of certain features in the sense of an emotion-induced confirmation bias, (2) certain features related to personal attitudes yielded arousal in individual participants, (3) this arousal affected the understanding of these variables.
@@ 1 Introduction and Related Work
More recently, the influence of emotions on the outcome of AI explanations has come into the focus of research. For example, it was shown that task-unrelated prior emotions affected advice taking behavior of participants given different explanation strategies [11]. More specifically, participants with high arousal were more likely to follow AI advice with a (guided) explanation than participants with low arousal. The latter preferred AI advice without any explanation.
On the other hand, explanations can induce affective reactions. Explanations given by an AI for an easy task were shown to yield negative effect in explainees whereas explanations in a difficult task caused positive affect. But also the quality of advice can affect the emotional state of an explainee. In a vignette study it was found that wrong advice would lead to negative feelings whereas correct advice would lead to positive feelings [2].
Interestingly, an investigation on the effectiveness of an XAI intervention showed that an explicit nudging strategy was successful in debiasing emotions [7]. However, it was not strong enough to debias decision making.
However, clear guidelines regarding how emotions affect the understanding in a decision-making situation and how an explaining system should take the explainee’s emotional state into account are still missing. This work contributes to the field by addressing this question.
@@ 2 Research Questions
In this research, we therefore investigate three research questions:
@@ 3.1.1 Independent Variables
Task-related emotions. We measure the emotional reactions by using the EmoNet [6] arousal data during the feature presentation and explanation. Based on the arousal value , an anomaly (emotional reaction) in the arousal state is detected by
@@ 3.1.3 Control Variables and Additional Measurements
mDES: [4] modified Differential Emotions Scale,
SAM [3]: Self-assessment manekin, providing a measure of participants’ self rated arousal and valence
@@
EmoNet [6] Results: In addition to arousal scores, full output vectors from the EmoNet model were stored for each facial frame, enabling detailed emotional state tracking over time.
@@ 3.2 Decision Suport System
In this study the same decision task and advice scheme was utilized as in a prior study [10]. The system centers around an embodied conversational agent named Flobi, who provides a personalized assessment of an individual’s risk profile based on a set of predefined input features. This risk assessment is subsequently used in the context of a Holt and Laury lottery task, where participants make incentivized decisions under risk. The agent’s evaluation serves as a form of decision support, offering guidance while allowing participants to ultimately make their own choices.
@@ 3.3 Participants
A total of 24 participants took part in the study. Of these, 15 were assigned to the fear condition and 9 to the happy condition. The sample included 8 male and 16 female participants. All participants were students at Bielefeld University.
@@ 3.4 Experimental Setting — Procedure
The experiment consisted of six phases (cf. Fig. 2. We recorded the users’ heart rate variability, video and audio, and their facial expressions as computed by EmoNet. In the first phase, a questionnaire was administered that requested data from the user to assess the user’s risk type. The risk type classification was achieved by a linear scoring scheme integrating the user’s numerical answers, yielding a risk type value between 0 and 9. After an emotion induction sequence, the actual risk task was explained to the user and the user could provide his/her risk selection, i.e., selecting a high or low risk on a scale from 0 to 9. After the user’s first risk decision, the system would present its risk suggestion to the user, based on the evaluation of the user’s risk type. More specifically, a user yielding a high value for a high-risk propensity would receive a suggestion of a high-risk choice and vice versa, also on a scale from 0 to 9. This suggestion was followed by an explanation of all ten variables and their relevance to the estimated risk type of the user. For example, being female was an indicator towards less risk propensity, whereas being male was an indicator for a high-risk propensity. We analyzed the HRV and Facial Expressions during theses episodes to detect arousal. After this explanation, the user could revise his/her decision. In the last phase, the users were asked to (verbally) name the features that they remembered from the explanation. After this, they were presented the features one after another and were asked to determine whether a certain value of this feature was an indicator for higher or lower risk propensity. We used this information as a proxy for understanding.
@@ 3.4.1 Questionnaire
During the first phase of the interaction, the participants were asked to fill out an online questionnaire, with the virtual robot Flobi being present and asking the questions. A total of ten questions were asked. We refer to these questions as the “variables” or “features” that the system uses to compute and explain the risk type of the participant.
Based on results from empirical studies reported in the literature for each variable, a scoring scheme was developed that assigned a score for each answer indicating higher or lower risk propensity. Overall, this resulted in a linear scoring scheme – the more scores, the higher the risk propensity. Scores were assigned for each feature based on a median value reported in the literature above or below which a higher (+1) or lower (0) risk score was given. This resulted in a risk type value between 0 and 9 for each participant.
@@ 3.4.4 XAI decision and explanation
In the explanation phase (phase 4, see above), Flobi would explain for each variable, whether the explainee’s (self-rated) value (e.g., of her political orientation) was an indicator adding to a higher or lower probability of the explainee as being more risk friendly. In this case, Flobi would say: “Because your political orientation is rather left, you are presumably less risk friendly”. Fig. 3 shows the GUI, where all variables are depicted with their respective contributions to the AI system’s decision. This would be repeated for all 11 variables.
@@ 3.4.5 Assessment of user’s retention and understanding of the explanations and the task
In the last phase, the uers’ retention and understanding were assessed using the procedures described in subsection 3.1.2. See Fig. 4 for the GUI to sort the features according to the remembered influence it had on the risk type classifcation.
@@ 4.1 Emotion Manipulation Check
To examine differences in affective responses between induction conditions, we compared SAM Valence and Arousal scores across groups (Fear vs. Happy - cf. Fig. 5). On the Valence scale, which ranges from 1 (“unpleasant feeling”) to 5 (“pleasant feeling”), participants in the Happy condition reported higher valence ratings (median 4), indicating more pleasant emotional states compared to the Fear group (median m 3). Similarly, on the Arousal scale (1 = “calm”, 5 = “aroused”), the Fear group tended to report higher arousal levels than the Happy group, suggesting that the fear induction elicited more physiologically activating emotional responses. These trends align with theoretical expectations – fear typically evokes high arousal and negative valence, while happiness is associated with positive valence and lower arousal.